In spite of the potential safety hazards, some people would like to have an Internet connection in their car. A preliminary survey of adult Americans has estimated this proportion to be somewhere around 0.30. Use the given preliminary estimate to determine the sample size required to estimate the proportion of adult Americans who would like an Internet connection in their car to within 0.02 with 90% confidence.
Answer #1

The sample size formula is given as below:

n = p*q*(Z/E)^2

We are given

p = 0.3

q = 1 – p = 0.7

Confidence level = 90%

Critical Z value = 1.6449

(by using z-table)

Margin of error = E = 0.02

The sample size is given as below:

n = p*q*(Z/E)^2

n = 0.3*0.7*(1.6449/0.02)^2

n = 1420.49

Required sample size = 1421